# arXiv:2603.02105v1[cs.NI] 02 Mar 2026

### v1

@@ Abstract
This paper presents the Distributed Adaptive Multi-Radio Cross-Layer Routing (DAMCR) protocol, designed to enhance reliability, adaptability, and energy efficiency in smart grid and industrial Internet of Things (IoT) communication networks. DAMCR integrates Chaotic Frequency-Hopping Spread Spectrum (C-FHSS) to improve physical-layer security and jamming resilience with Link-Adaptive Quality Power Control (LAQPC) to dynamically regulate transmission power based on instantaneous link quality and residual node energy. To meet heterogeneous traffic requirements, the protocol incorporates priority-aware message classification that differentiates between periodic monitoring data and time-critical fault and protection messages. The proposed framework is implemented and evaluated in MATLAB using a heterogeneous network composed of LoRa, Wi-Fi, and dual-radio nodes operating under AWGN, Rayleigh, and Rician fading environments. Extensive simulation results demonstrate that DAMCR consistently achieves a Packet Delivery Ratio (PDR) exceeding 95% across all evaluated scenarios, while maintaining end-to-end latency between 17 and 23 ms, even in the presence of controlled jamming attacks. These results confirm that the tight integration of chaos-based spectrum agility, cross-technology routing, and energy-aware cross-layer adaptation significantly improves communication reliability, latency stability, and resilience compared to conventional single-radio and static-routing protocols.
@@ VI Conclusion
Future research will extend DAMCR toward real-world deployment using hybrid LoRa–Wi-Fi hardware or software-defined radio platforms to validate real-time synchronisation and energy performance. Incorporating mobility-aware routing and cognitive spectrum sensing could further enhance adaptability in dynamic or congested environments. In addition, integrating energy-harvesting capabilities and lightweight cryptographic schemes alongside chaotic modulation may enable sustainable, self-securing IoT networks. The proposed framework, therefore, establishes a foundational step toward resilient, secure, and energy-aware communication architectures for next-generation smart grid and industrial IoT systems.

### v2

# Resilient Chaotic Cross-Layer Routing for Smart Grid IoT Networks

@@ Abstract
This paper presents the Distributed Adaptive Multi-Radio Cross-Layer Routing (DAMCR) protocol, designed to enhance reliability, adaptability, and energy efficiency in smart grid and industrial Internet of Things (IoT) communication networks. DAMCR integrates Chaotic Frequency-Hopping Spread Spectrum (C-FHSS) to improve physical-layer security and jamming resilience with Link-Adaptive Quality Power Control (LAQPC) to dynamically regulate transmission power based on instantaneous link quality and residual node energy. To meet heterogeneous traffic requirements, the protocol incorporates priority-aware message classification that differentiates between periodic monitoring data and time-critical fault and protection messages. The proposed framework is implemented and evaluated in MATLAB using a heterogeneous network composed of LoRa, Wi-Fi, and dual-radio nodes operating under AWGN, Rayleigh, and Rician fading environments. Extensive simulation results demonstrate that DAMCR consistently achieves a Packet Delivery Ratio (PDR) exceeding 98% across all evaluated scenarios, while maintaining end-to-end latency between 17 and 23 ms, even under controlled jamming attacks. These results confirm that the tight integration of chaos-based spectrum agility, cross-technology routing, and energy-aware cross-layer adaptation significantly improves communication reliability, latency stability, and resilience compared to conventional single-radio and static-routing protocols.
@@ VI Conclusion
Future research will extend DAMCR toward real-world deployment using hybrid LoRa–Wi-Fi hardware or software-defined radio platforms to validate real-time synchronisation and energy performance. Incorporating mobility-aware routing and cognitive spectrum sensing could further enhance adaptability in dynamic or congested environments. Integrating energy-harvesting capabilities and lightweight cryptographic schemes with chaotic modulation may enable sustainable, self-securing IoT networks, laying the foundation for resilient, secure, and energy-aware communication architectures for next-generation smart grid and industrial IoT systems.